A random sample of 722 residents in a major town was asked whether they had ever been bitten by a dog. The responses (1=Yes and 2=N0) are recorded. Estimate with 95% confidence the proportion of residents who have been bitten by a dog.

Sample frequencies: n (1) = 304; n (2) = 418


1
Expert's answer
2021-03-17T12:07:14-0400

The sample proportion is computed as follows, based on the sample size "N=722"  and the number of favorable cases "X=304"


"\hat{p}=\dfrac{X}{N}=\dfrac{304}{722}\approx 0.42105"

The critical value for "\alpha=0.05" is "z_c=z_{1-\alpha/2}=1.96."

The corresponding confidence interval is computed as shown below:


"CI(Proportion)"

"=(\hat{p}-z_c\sqrt{\dfrac{\hat{p}(1-\hat{p})}{n}},\hat{p}+z_c\sqrt{\dfrac{\hat{p}(1-\hat{p})}{n}})"

"=(0.42105-1.96\sqrt{\dfrac{0.42105(1-0.42105)}{722}},"

"0.42105+1.96\sqrt{\dfrac{0.42105(1-0.42105)}{722}})"


"=(0.3850, 0.4571)"

Therefore, based on the data provided, the 95% confidence interval for the population proportion is "0.3850<p<0.4571," which indicates that we are 95% confident that the true population proportion "p" is contained by the interval "(0.3850, 0.4571)."